1.6.5. f(z) = ?? 1.6.6. f(z) = sin. 1.6.7. f(3) = 3 1.6.8. f(z) = cos 2z.


For 1.6.5-8, verify the Cauchy-Riemann equations for the functions.
Answer #1

riemann equation is sahisted cauchy - t het pro eo en e me Dva -sin(20) coshczy) -2005(2x)sinhley) ke ne be = ne (Re)44 15 (x​​​​​​. -1795 5.91 f(2)= 23h 0 let f (2) = 4(49) + iev(x, y) ze atij . f(z) = (ariy;} => 418,9)+1¢v(x,y) = (23-347?) ti( 34%y-y3) . 7» Buong by cauchy ziemon satisfied equation is (021618 ] ] Recept he movement he alghtak) (he) hat (sht qe)2 = lahtel The tak